OpenStudy (afloridagirl):

Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? Left by 4 units Right by 4 units Left by 8 units Right by 8 units

OpenStudy (afloridagirl):

@jim_thompson5910

OpenStudy (afloridagirl):

i thought it was a but it was wrong

jimthompson5910 (jim_thompson5910):

x^2+8x+16 factors to what?

OpenStudy (afloridagirl):

im not sure to tell the truth

jimthompson5910 (jim_thompson5910):

think of ways to multiply to 16 16 = 1*16 16 = 2*8 16 = 4*4 which of those pairs of factors add up to 8? is it 1 and 16? 2 and 8? 4 and 4?

OpenStudy (afloridagirl):

4 and 4?

jimthompson5910 (jim_thompson5910):

so x^2+8x+16 factors to (x+4)(x+4) which is the same as (x+4)^2

OpenStudy (afloridagirl):

right so that means that it would be b?

jimthompson5910 (jim_thompson5910):

To go from x^2 to (x+4)^2, we shift 4 units to the left so choice A is correct. Your teacher messed up

jimthompson5910 (jim_thompson5910):

if you wanted to shift 4 units right,you'd go from x^2 to (x-4)^2

OpenStudy (afloridagirl):

???

OpenStudy (afloridagirl):

so not b

jimthompson5910 (jim_thompson5910):

https://www.desmos.com/calculator/2wcevwcxik

jimthompson5910 (jim_thompson5910):

Notice how if we shifted the red graph 4 units to the left, we would end up at the blue graph

OpenStudy (afloridagirl):

yes i believe

jimthompson5910 (jim_thompson5910):

The answer should be A. Your teacher made a typo somewhere

OpenStudy (afloridagirl):

oh okay thank you for your help!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (afloridagirl):

it still says that that one is wrong....

jimthompson5910 (jim_thompson5910):

You'll have to discuss it with your teacher. S/he will be able to fix it